OpenStudy (anonymous):

Find the roots of x^3+2x^2-5x-6=0

OpenStudy (anonymous):

just factor (x-2)*(x+3)*(x+1)=0 can you solve that?

OpenStudy (anonymous):

Would you do the separatly? I do not get how that works.

OpenStudy (anonymous):

once you factor you can set each individual ( ) to 0 ( )=0 you should get 3 roots

OpenStudy (anonymous):

x-2=0 x+3=0 x+1= 0 solve each one

OpenStudy (anonymous):

So would it be 2,-3,-1?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

if you have a graphing calculator you can graph it and find where x intersect 0